OpenStudy (anonymous):

let r be the region in the first quadrant bounded by the graph y=8- x^ (3/2) Find the area of the region R . Find the volume of the solid generated when R is revolved about the x-axis

OpenStudy (amistre64):

you sure its x^(3/2) ?

OpenStudy (amistre64):

the function itself is of no consequence; it only helps define the radial limit and the boundary for integration

OpenStudy (amistre64):

when does y = 0?

OpenStudy (anonymous):

yes its x^3/2

OpenStudy (anonymous):

i typed allthe given info

OpenStudy (amistre64):

ok, cause the last question i just did on this was an x^2 there :)

OpenStudy (amistre64):

yes, and all the given info allows us to find the needed info

OpenStudy (amistre64):

when does y=0?

OpenStudy (amistre64):



OpenStudy (anonymous):

i'm so lost

OpenStudy (amistre64):

\[\int_{y=8}^{y=0}(area.of.a.circle)dx\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (amistre64):

y = 8-x^(3/2); when does y=0? 0 = 8-x^(3/2) ; solve for x

OpenStudy (amistre64):

thats a guide; thats what we are going to have to fill in to answer the question

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

o wait no

OpenStudy (amistre64):

well, if were gonna be guessing; lets at least fill it in to see :)

OpenStudy (anonymous):

x=4?

OpenStudy (amistre64):

4 is better

OpenStudy (anonymous):

i forgot to square it, my bad

OpenStudy (amistre64):

\[\int_{0}^{4}(area.function)dx\]

OpenStudy (amistre64):

each rotation produces a circle; what the formula for the area of a circle?

OpenStudy (anonymous):

pi r sqd

OpenStudy (amistre64):

good; except for in this case; r = f(x) as we move down the line

OpenStudy (anonymous):

so how do we get r

OpenStudy (amistre64):

\[\int_{0}^{4}\pi[r]^2dx\] \[\int_{0}^{4}\pi[f(x)]^2dx\]

OpenStudy (anonymous):

f(x)=y

OpenStudy (anonymous):

subsitute 4 for x?

OpenStudy (amistre64):